OpenStudy (anonymous):

Complete the square for the expression. Write the resulting expression as a binomial squared. x2 - 9x + _____

OpenStudy (anonymous):

(x-4,5)^2 will give you x^2 - 9x + 81/4 (or 20.25, depending on how you want to write it)

OpenStudy (anonymous):

ax²+bx+c To complete the square (find c) divide b by 2 and square that number. x² - 9x + ____? (9/2)² = 81/4 = x² - 9x + 81/4 = (x-9/2)²